ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 194, 2013

Asaba, Y. and A. Hanaki A construction of integral standard generalized table algebras from parameters of projective geometries

395

Bačák, M. The proximal point algorithm in metric spaces  689

Balogh, J. and J. Lenz On the Ramsey–Turán numbers of graphs and hypergraphs

45

Bartholdi, L. See A. Smoktunowicz Baruch, E. M., E. Lapid and Z. Mao A Bessel identity for the

theta correspondence

225

Bary-Soroker, L. and E. Paran Fully Hilbertian fields  507
Ben Yaacov, I. On theories of random variables  957

Berkovich, Y. Maximal abelian and minimal nonabelian subgroups of some finite two-generator p-groups especially metacyclic

831

Bourgain, J. Prescribing the binary digits of primes  935

Bowen, L. and R. D. Tucker-Drob On a co-induction question of Kechris

209

Chan, H. H., J. Wan and W. Zudilin Legendre polynomials and Ramanujan-type series for 1/π

183

Chernikov, A. and P. Simon Externally definable sets and dependent pairs

409

Chorny, B. Brown representability for space-valued functors  767

Colombo, F., I. Sabadini and F. Sommen The inverse Fueter mapping theorem in integral form using spherical monogenics

485

Costello, K. P. Bilinear and quadratic variants on the Littlewood– Offord problem

359

Couveignes, J.-M. and R. Lercier Fast construction of irreducible polynomials over finite fields

77

Frantzikinakis, N., B. Host and B. Kra The polynomial multidimensional Szemerédi Theorem along shifted primes

331

Ghioca, D. and T. Scanlon Algebraic equations on the adèlic closure of a Drinfeld module

461

Gimeno, V. and V. Palmer Extrinsic isoperimetry and compactification of minimal surfaces in Euclidean and hyperbolic spaces

539

Grbić, J., J. Harper, M. Mimura, S. Theriault and J. Wu Rank p − 1 mod-p H-spaces 641 Green, B. Bounds on the number of automorphisms of curves over

algebraically closed fields

69

Hanaki, A. See Y. Asaba Harper, J. See J. Grbić Host, B. See N. Frantzikinakis

Juhász, I., J. van Mill and W. Weiss Variations on ω-boundedness  745
Kalton, N. J. Examples of uniformly homeomorphic Banach spaces  151

Keller, M., D. Lenz and S. Warzel On the spectral theory of trees with finite cone type

107

Kra, B. See N. Frantzikinakis

Krupiński, K. and F. Wagner Small, nm-stable compact G-groups  907
Lanconelli, E. “Potato kugel” for sub-Laplacians  277

Lapid, E. See E. M. Baruch Lenz, D. See M. Keller Lenz, J. See J. Balogh Lercier, R. See J.-M. Couveignes Liskevich, V. and I. I. Skrypnik Pointwise estimates for solutions

to the porous medium equation with measure as a forcing term

259

Liu, X. Y. and Z. H. Liu Existence results for a class of second order evolution inclusions and its corresponding first order evolution inclusions

723

Liu, Z. H. See X. Y. Liu

Lorscheid, O. Toroidal automorphic forms for function fields  555

Mao, Z. See E. M. Baruch Melbourne, I. and D. Terhesiu First and higher order uniform

dual ergodic theorems for dynamical systems with infinite measure

793

van Mill, J. See I. Juhász Mimura, M. See J. Grbić Palmer, V. See V. Gimeno Paran, E. See L. Bary-Soroker

Reeve, H. WJ. Infinite non-conformal iterated function systems  285

Repovš, D. and M. Zaicev On the codimension growth of almost nilpotent Lie algebras

137

Sabadini, I. See F. Colombo Scanlon, T. See D. Ghioca

Serafin, J. Non-existence of a universal zero-entropy system  349

Shumyatsky, P. Exponent of a finite group with a fixed-point-free four-group of automorphisms

895

Simon, P. See A. Chernikov Skrypnik, I. I. See V. Liskevich Smoktunowicz, A. and L. Bartholdi Jacobson radical non-nil al-

gebras of Gel′fand–Kirillov dimension 2

597

Soifer, G. A. Affine semigroups acting properly discontinuously on a hyperbolic space

703

Sokić, M. Ramsey property, ultrametric spaces, finite posets, and universal minimal flows

609

Solanes, G. and E. Teufel Horo-tightness and total (absolute) curvatures in hyperbolic spaces

427

Sommen, F. See F. Colombo Terhesiu, D. See I. Melbourne Teufel, E. See G. Solanes Theriault, S. See J. Grbić Tucker-Drob, R. D. See L. Bowen Venketasubramanian, C. G. On representations of GL(n) distin-

guished by GL(n − 1) over a p-adic field

1

Wagner, F. See K. Krupiński Wan, J. See H. H. Chan Warzel, S. See M. Keller

Wei, J. Semi-tilting complexes  871

Weiss, W. See I. Juhász Wu, J. See J. Grbić Zaicev, M. See D. Repovš Zudilin, W. See H. H. Chan